OpenStudy (anonymous):

radical 10 times radical 8

OpenStudy (anonymous):

\(\large \sqrt{10} \cdot \sqrt{8}=\sqrt{2} \cdot\sqrt{5} \cdot\sqrt{2} \cdot \sqrt{4} \) = \(\large \sqrt{2}\cdot\sqrt{2} \cdot \sqrt{4} \cdot\sqrt{5} \) = \(\large 2 \cdot 2 \cdot\sqrt{5} \) = \(\large 4\sqrt5 \)

OpenStudy (anonymous):

How do you type the radical sign? @dpaInc

OpenStudy (anonymous):

you'll need to enter it as LaTex... right click the text you want to see to see the code.... choose "Show Math As" ---> "Tex commands"

OpenStudy (anonymous):

or you can just use that blue "Equation" button at the lower left of this text box...

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw... :)